OpenStudy (anonymous):

Subtract... 8/10 - 1/4 PLEASE HELP ME

OpenStudy (anonymous):

PLEASE I am DESPRATE

OpenStudy (anonymous):

.55 = 11/20

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

1st find a common denominator 2nd subtract 3rd figure out common denominator is 40 so then subtract and then u get 22/40 then simplified = 11/20

OpenStudy (anonymous):

Find the common denominator, subtract and presto...you have an answer.